OpenStudy (anonymous):

how do you find the solution set for 2x^2 is greater than or equal to 8?

OpenStudy (nincompoop):

have you tried 2?

OpenStudy (anonymous):

Does it looks like \[2x^\ge 8\]

OpenStudy (nincompoop):

NO

OpenStudy (anonymous):

First replace inequality sign by equality so we get 2x^2 = 8

OpenStudy (nincompoop):

you don't need to

OpenStudy (nincompoop):

\[\frac{ 2x^2 }{ 2 } \ge \frac{ 8 }{ 2 } \ge \sqrt{x^2} \ge \sqrt{4}\]

OpenStudy (nincompoop):

solve for x by merely isolating it

OpenStudy (nincompoop):

get rid of 2 by dividing both sides by 2 then get rid of the exponent by obtaining it's equivalent root - the square root

OpenStudy (anonymous):

I am lost

OpenStudy (anonymous):

I know 2 is the square root of 4 but how does that give you the solution set?